# Constrained Stochastic Recursive Momentum Successive Convex Approximation

### v3

@@ VI Conclusion
We have considered the problem of minimizing a stochastic non-convex optimization problem with non-convex constraints. While the problem has been well-studied, the state-of-the-art algorithms have suboptimal oracle complexity of ${\mathcal{O}}(\epsilon^{-2})$. The recently proposed algorithm in [14] achieves the optimal rate of ${\mathcal{O}}(\epsilon^{-3/2})$, but this result is only valid under the condition that the initial point is nearly feasible with non-adaptive step sizes. In contrast, we propose the CoSTA algorithm, which introduces recursive momentum updates to achieve an almost optimal oracle complexity of ${\tilde{\mathcal{O}}}(\epsilon^{-3/2})$ under adaptive step size, and achieves an optimal oracle complexity with non-adaptive step size. The proposed algorithm is developed within the SCA framework, and requires convex surrogates of the objective and constraint functions at every iteration. We also introduced a parameterized version of the Mangasarian-Fromovitz constraint qualification that allows us to ensure that the intermediate subproblems are strictly feasible and the intermediate dual variable stay bounded, resulting in an oracle complexity bound that depends only on the problem parameters. Finally, the empirical performance of the proposed CoSTA algorithm is compared with the other state-of-the-art algorithms across two different applications: energy optimal trajectory design and sparse binary classification. The proposed algorithm outperforms the state-of-the-art algorithms on the energy optimal trajectory design problem and performs at par with the LCPP algorithm on the sparse binary classification problem.
@@ Appendix A Proof of Lemma 1
Proof:
[…]
${\mathbb{E}}_{{\boldsymbol{\xi}}_{t+1}}[\tfrac{\left\|{\mathbf{z}}_{t+2}-\nabla U({\mathbf{x}}_{t+1})\right\|^{2}}{\eta_{t}}]$ $={\mathbb{E}}_{{\boldsymbol{\xi}}_{t+1}}\bigg{[}\frac{(1-\beta_{t+1})^{2}\left\|{\mathbf{z}}_{t+1}-\nabla U({\mathbf{x}}_{t})\right\|^{2}}{\eta_{t}}\bigg{]}+{\mathbb{E}}_{{\boldsymbol{\xi}}_{t+1}}\Big{[}\eta_{t}^{-1}\big{\|}(1-\beta_{t+1})\big{(}\nabla f({\mathbf{x}}_{t},{\boldsymbol{\xi}}_{t+1})$
[…]
${\mathbb{E}}_{{\boldsymbol{\xi}}_{t+1}}\Big{[}\eta_{t}^{-1}\big{\|}(1-\beta_{t+1})\big{(}\nabla f({\mathbf{x}}_{t},{\boldsymbol{\xi}}_{t+1})-\nabla U({\mathbf{x}}_{t})+\nabla U({\mathbf{x}}_{t+1})-\nabla f({\mathbf{x}}_{t+1},{\boldsymbol{\xi}}_{t+1})\big{)}$
$~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}-\beta_{t+1}\big{(}\nabla f({\mathbf{x}}_{t+1},{\boldsymbol{\xi}}_{t+1})-\nabla U({\mathbf{x}}_{t+1}))\big{\|}^{2}\Big{]}$
$\overset{(a)}{\leq}2{\mathbb{E}}_{{\boldsymbol{\xi}}_{t+1}}\bigg{[}\frac{(1-\beta_{t+1})^{2}\left\|\nabla f({\mathbf{x}}_{t+1},{\boldsymbol{\xi}}_{t+1})-\nabla f({\mathbf{x}}_{t},{\boldsymbol{\xi}}_{t+1})\right\|^{2}}{\eta_{t}}\bigg{]}+2{\mathbb{E}}_{{\boldsymbol{\xi}}_{t+1}}\bigg{[}\frac{\beta_{t+1}^{2}\left\|\nabla f({\mathbf{x}}_{t+1},{\boldsymbol{\xi}}_{t+1})-\nabla U({\mathbf{x}}_{t+1})\right\|^{2}}{\eta_{t}}\bigg{]}$
[…]
Using $L$-smoothness, and the update equation (10), we obtain
[…]
$\leq\frac{1}{16L^{2}}{\mathbb{E}}\left[\frac{(1-\beta_{t+1})\left\|{{\boldsymbol{\varepsilon}}}_{t}\right\|^{2}}{\eta_{t}}\right]+\frac{1}{8}{\mathbb{E}}\left[\eta_{t}\left\|{\boldsymbol{\delta}}_{t}\right\|^{2}\right]+\frac{1}{8L^{2}}{\mathbb{E}}\left[\frac{\beta_{t+1}^{2}G_{t+1}^{2}}{\eta_{t}}\right].$ (79)
Appendix B Proof of Lemma 2
Proof:
$\langle{{\hat{{\mathbf{x}}}}_{t}}-{\mathbf{x}}_{t},{\mathbf{z}}_{t+1}+{\mathbf{v}}_{{\hat{{\mathbf{x}}}}_{t}}\rangle$ $~{}+\langle{{\hat{{\mathbf{x}}}}_{t}}-{\mathbf{x}}_{t},\nabla{\tilde{f}}({{\hat{{\mathbf{x}}}}_{t}},{\mathbf{x}}_{t},{\mathbf{z}}_{t},{\boldsymbol{\xi}}_{t})-\nabla{\tilde{f}}({\mathbf{x}}_{t},{\mathbf{x}}_{t},{\mathbf{z}}_{t},{\boldsymbol{\xi}}_{t})\rangle\leq 0.$ (82)
$\Rightarrow\langle{\mathbf{z}}_{t+1}+{\mathbf{v}}_{{\hat{{\mathbf{x}}}}_{t}},{{\hat{{\mathbf{x}}}}_{t}}-{\mathbf{x}}_{t}\rangle+\mu\left\|{{\hat{{\mathbf{x}}}}_{t}}-{\mathbf{x}}_{t}\right\|^{2}\leq 0.$ (83)
Multiplying (83) by $\eta_{t}$ and adding, we obtain
Adding and subtracting $\nabla U({\mathbf{x}}_{t})$, we obtain
$\eta_{t}\langle{\mathbf{z}}_{t+1}-\nabla U({\mathbf{x}}_{t}),{{\hat{{\mathbf{x}}}}_{t}}-{\mathbf{x}}_{t}\rangle$ $+\eta_{t}\langle\nabla U({\mathbf{x}}_{t}),{{\hat{{\mathbf{x}}}}_{t}}-{\mathbf{x}}_{t}\rangle$
$~{}~{}~{}\leq-\mu\eta_{t}\left\|{{\hat{{\mathbf{x}}}}_{t}}-{\mathbf{x}}_{t}\right\|^{2}+u({\mathbf{x}}_{t})-u({\mathbf{x}}_{t+1})$ (86)
$\leq U({\mathbf{x}}_{t})+u({\mathbf{x}}_{t})-u({\mathbf{x}}_{t+1})-\eta_{t}\langle{\mathbf{z}}_{t+1}-\nabla U({\mathbf{x}}_{t}),{{\hat{{\mathbf{x}}}}_{t}}-{\mathbf{x}}_{t}\rangle+\bigg{(}\frac{L}{2}\eta_{t}^{2}-\mu\eta_{t}\bigg{)}\left\|{{\hat{{\mathbf{x}}}}_{t}}-{\mathbf{x}}_{t}\right\|^{2}$ (89)
${\mathbb{E}}\left[F({\mathbf{x}}_{t+1})-F({\mathbf{x}}_{t})\right]$ $\leq\frac{1}{2}{\mathbb{E}}\left[\eta_{t}\left\|\nabla U({\mathbf{x}}_{t})-{\mathbf{z}}_{t+1}\right\|^{2}\right]+\frac{L}{2}{\mathbb{E}}\left[\eta_{t}^{2}\left\|{{\hat{{\mathbf{x}}}}_{t}}-{\mathbf{x}}_{t}\right\|^{2}\right]+\bigg{(}\frac{1}{2}-\mu\bigg{)}{\mathbb{E}}\left[\eta_{t}\left\|{{\hat{{\mathbf{x}}}}_{t}}-{\mathbf{x}}_{t}\right\|^{2}\right]$
Appendix C Proof of Lemma 4
@@
$={\tilde{g}}_{j}({\mathbf{x}}_{t})+\gamma\langle\nabla{\tilde{g}}_{j}({\mathbf{x}}_{t}),{\mathbf{d}}_{t}\rangle+\frac{\gamma^{2}L}{2}\leq-\rho\gamma+\frac{\gamma^{2}L}{2}$ (92)
@@
Appendix D Proof of Lemma 5
@@
${\tilde{f}}({\mathbf{x}},{\mathbf{x}}_{t},{\mathbf{z}}_{t},{\boldsymbol{\xi}}_{t})7+u({\mathbf{x}})$ $\geq{\tilde{f}}({{\hat{{\mathbf{x}}}}_{t}},{\mathbf{x}}_{t},{\mathbf{z}}_{t},{\boldsymbol{\xi}}_{t})+u({{\hat{{\mathbf{x}}}}_{t}})+\langle\nabla{\tilde{f}}({{\hat{{\mathbf{x}}}}_{t}},{\mathbf{x}}_{t},{\mathbf{z}}_{t},{\boldsymbol{\xi}}_{t})+{\mathbf{v}}_{{{\hat{{\mathbf{x}}}}_{t}}},{\mathbf{x}}-{{\hat{{\mathbf{x}}}}_{t}}\rangle$ (97)
@@
${\tilde{f}}({\tilde{\mathbf{x}}}({\mathbf{x}}_{t}),{\mathbf{x}}_{t},{\mathbf{z}}_{t},{\boldsymbol{\xi}}_{t})+u({\tilde{\mathbf{x}}}({\mathbf{x}}_{t}))+\sum\nolimits_{j=1}^{J}[{\hat{{\boldsymbol{\lambda}}}}_{t}]_{j}{\tilde{g}}_{j}({\tilde{\mathbf{x}}}({\mathbf{x}}_{t}),{\mathbf{x}}_{t})+\sum\nolimits_{i=1}^{I}[\hat{\nu}_{t}]_{i}h_{i}({\tilde{\mathbf{x}}}({\mathbf{x}}_{t}))$
$\geq{\tilde{f}}({{\hat{{\mathbf{x}}}}_{t}},{\mathbf{x}}_{t},{\mathbf{z}}_{t},{\boldsymbol{\xi}}_{t})+u({{\hat{{\mathbf{x}}}}_{t}})+\sum\nolimits_{j=1}^{J}[{\hat{{\boldsymbol{\lambda}}}}_{t}]_{j}{\tilde{g}}_{j}({{\hat{{\mathbf{x}}}}_{t}},{\mathbf{x}}_{t})+\sum\nolimits_{i=1}^{I}[\hat{\nu}_{t}]_{i}h_{i}({{\hat{{\mathbf{x}}}}_{t}})$
$+\Big{\langle}\sum\nolimits_{j=1}^{J}[{\hat{{\boldsymbol{\lambda}}}}_{t}]_{j}\nabla{\tilde{g}}_{j}({{\hat{{\mathbf{x}}}}_{t}},{\mathbf{x}}_{t})+\sum\nolimits_{i=1}^{I}[\hat{\nu}_{t}]_{i}\nabla h_{i}({{\hat{{\mathbf{x}}}}_{t}})+\nabla{\tilde{f}}({{\hat{{\mathbf{x}}}}_{t}},{\mathbf{x}}_{t},{\mathbf{z}}_{t},{\boldsymbol{\xi}}_{t})+{\mathbf{v}}_{{{\hat{{\mathbf{x}}}}_{t}}},{\tilde{\mathbf{x}}}({\mathbf{x}}_{t})-{{\hat{{\mathbf{x}}}}_{t}}\Big{\rangle}$ (100)

### v4

@@ VI Conclusion
We have considered the problem of minimizing a stochastic non-convex optimization problem with non-convex constraints. While the problem has been well-studied, the state-of-the-art algorithms have suboptimal oracle complexity of ${\mathcal{O}}(\epsilon^{-2})$. The recently proposed algorithm in [14] achieves the optimal rate of ${\mathcal{O}}(\epsilon^{-3/2})$, but this result is only valid under the condition that the initial point is nearly feasible with non-adaptive step sizes. In contrast, we propose the CoSTA algorithm, which introduces recursive momentum updates to achieve an almost optimal oracle complexity of ${\tilde{\mathcal{O}}}(\epsilon^{-3/2})$ under adaptive step sizes and Lipschitz smoothness assumptions. Furthermore, under a milder Mean Squared Smoothness (MSS) condition, CoSTA retains this complexity with non-adaptive step sizes. The proposed algorithm is developed within the SCA framework, and requires convex surrogates of the objective and constraint functions at every iteration. We also introduced a parameterized version of the Mangasarian-Fromovitz constraint qualification that allows us to ensure that the intermediate subproblems are strictly feasible and the intermediate dual variable stay bounded, resulting in an oracle complexity bound that depends only on the problem parameters. Finally, the empirical performance of the CoSTA algorithm is compared with the other state-of-the-art algorithms across two different applications: energy optimal trajectory design and sparse binary classification. The proposed algorithm outperforms the state-of-the-art algorithms on the energy optimal trajectory design problem and performs at par with the LCPP algorithm on the sparse binary classification problem.
@@ Appendix A Proof of Lemma 4
$\hskip-42.67912pt={\tilde{g}}_{j}({\mathbf{x}}_{t})+\gamma\langle\nabla{\tilde{g}}_{j}({\mathbf{x}}_{t}),{\mathbf{d}}_{t}\rangle+\frac{\gamma^{2}L}{2}\leq-\rho\gamma+\frac{\gamma^{2}L}{2}$ (74)
@@
Appendix B Proof of Lemma 5
@@
${\tilde{f}}({\mathbf{x}},{\mathbf{x}}_{t},{\mathbf{z}}_{t},{\boldsymbol{\xi}}_{t})$ $+u({\mathbf{x}})\geq{\tilde{f}}({{\hat{{\mathbf{x}}}}_{t}},{\mathbf{x}}_{t},{\mathbf{z}}_{t},{\boldsymbol{\xi}}_{t})+u({{\hat{{\mathbf{x}}}}_{t}})$
$+\langle\nabla{\tilde{f}}({{\hat{{\mathbf{x}}}}_{t}},{\mathbf{x}}_{t},{\mathbf{z}}_{t},{\boldsymbol{\xi}}_{t})+{\mathbf{v}}_{{{\hat{{\mathbf{x}}}}_{t}}},{\mathbf{x}}-{{\hat{{\mathbf{x}}}}_{t}}\rangle$ (79)
@@
${\tilde{f}}({\tilde{\mathbf{x}}}({\mathbf{x}}_{t}),{\mathbf{x}}_{t},{\mathbf{z}}_{t},{\boldsymbol{\xi}}_{t})+u({\tilde{\mathbf{x}}}({\mathbf{x}}_{t}))$
$+\sum\nolimits_{j=1}^{J}[{\hat{{\boldsymbol{\lambda}}}}_{t}]_{j}{\tilde{g}}_{j}({\tilde{\mathbf{x}}}({\mathbf{x}}_{t}),{\mathbf{x}}_{t})+\sum\nolimits_{i=1}^{I}[\hat{\nu}_{t}]_{i}h_{i}({\tilde{\mathbf{x}}}({\mathbf{x}}_{t}))$
$\geq{\tilde{f}}({{\hat{{\mathbf{x}}}}_{t}},{\mathbf{x}}_{t},{\mathbf{z}}_{t},{\boldsymbol{\xi}}_{t})+u({{\hat{{\mathbf{x}}}}_{t}})$
$~~~+\sum\nolimits_{j=1}^{J}[{\hat{{\boldsymbol{\lambda}}}}_{t}]_{j}{\tilde{g}}_{j}({{\hat{{\mathbf{x}}}}_{t}},{\mathbf{x}}_{t})+\sum\nolimits_{i=1}^{I}[\hat{\nu}_{t}]_{i}h_{i}({{\hat{{\mathbf{x}}}}_{t}})$
$+\Big{\langle}\sum\nolimits_{j=1}^{J}[{\hat{{\boldsymbol{\lambda}}}}_{t}]_{j}\nabla{\tilde{g}}_{j}({{\hat{{\mathbf{x}}}}_{t}},{\mathbf{x}}_{t})+\sum\nolimits_{i=1}^{I}[\hat{\nu}_{t}]_{i}\nabla h_{i}({{\hat{{\mathbf{x}}}}_{t}})$
$~~~+\nabla{\tilde{f}}({{\hat{{\mathbf{x}}}}_{t}},{\mathbf{x}}_{t},{\mathbf{z}}_{t},{\boldsymbol{\xi}}_{t})+{\mathbf{v}}_{{{\hat{{\mathbf{x}}}}_{t}}},{\tilde{\mathbf{x}}}({\mathbf{x}}_{t})-{{\hat{{\mathbf{x}}}}_{t}}\Big{\rangle}$ (82)
@@
Appendix C Proof of Lemma 1
Proof:
${\mathbf{z}}_{t+2}-$ $\nabla U({\mathbf{x}}_{t+1})=(1-\beta_{t+1})({\mathbf{z}}_{t+1} -\nabla U({\mathbf{x}}_{t}))$
$-(1-\beta_{t+1})(\nabla f({\mathbf{x}}_{t},{\boldsymbol{\xi}}_{t+1})-\nabla U({\mathbf{x}}_{t}))$
${\mathbb{E}}_{{\boldsymbol{\xi}}_{t+1}}[\tfrac{\left\|{\mathbf{z}}_{t+2}-\nabla U({\mathbf{x}}_{t+1})\right\|^{2}}{\eta_{t}}]$
$~={\mathbb{E}}_{{\boldsymbol{\xi}}_{t+1}}\bigg{[}\frac{(1-\beta_{t+1})^{2}\left\|{\mathbf{z}}_{t+1}-\nabla U({\mathbf{x}}_{t})\right\|^{2}}{\eta_{t}}\bigg{]}$
$~~~~+{\mathbb{E}}_{{\boldsymbol{\xi}}_{t+1}}\Big{[}\eta_{t}^{-1}\big{\|}(1-\beta_{t+1})\big{(}\nabla f({\mathbf{x}}_{t},{\boldsymbol{\xi}}_{t+1})$
$~~~~~~~~-\nabla U({\mathbf{x}}_{t})+\nabla U({\mathbf{x}}_{t+1})-\nabla f({\mathbf{x}}_{t+1},{\boldsymbol{\xi}}_{t+1})\big{)}$
[…]
${\mathbb{E}}_{{\boldsymbol{\xi}}_{t+1}}\Big{[}\eta_{t}^{-1}\big{\|}(1-\beta_{t+1})\big{(}\nabla f({\mathbf{x}}_{t},{\boldsymbol{\xi}}_{t+1})$
$~~~~~-\nabla U({\mathbf{x}}_{t})+\nabla U({\mathbf{x}}_{t+1})-\nabla f({\mathbf{x}}_{t+1},{\boldsymbol{\xi}}_{t+1})\big{)}$
$~~~~~~~~~~~-\beta_{t+1}\big{(}\nabla f({\mathbf{x}}_{t+1},{\boldsymbol{\xi}}_{t+1})-\nabla U({\mathbf{x}}_{t+1}))\big{\|}^{2}\Big{]}$
$\overset{(a)}{\leq}2{\mathbb{E}}_{{\boldsymbol{\xi}}_{t+1}}\bigg{[}\frac{(1-\beta_{t+1})^{2}\left\|\nabla f({\mathbf{x}}_{t+1},{\boldsymbol{\xi}}_{t+1})-\nabla f({\mathbf{x}}_{t},{\boldsymbol{\xi}}_{t+1})\right\|^{2}}{\eta_{t}}\bigg{]}$
$~~~~~~~~~~+2{\mathbb{E}}_{{\boldsymbol{\xi}}_{t+1}}\bigg{[}\frac{\beta_{t+1}^{2}\left\|\nabla f({\mathbf{x}}_{t+1},{\boldsymbol{\xi}}_{t+1})-\nabla U({\mathbf{x}}_{t+1})\right\|^{2}}{\eta_{t}}\bigg{]}$
$\overset{(b)}{\leq}2{\mathbb{E}}_{{\boldsymbol{\xi}}_{t+1}}\bigg{[}\frac{(1-\beta_{t+1})^{2}\left\|\nabla f({\mathbf{x}}_{t+1},{\boldsymbol{\xi}}_{t+1})-\nabla f({\mathbf{x}}_{t},{\boldsymbol{\xi}}_{t+1})\right\|^{2}}{\eta_{t}}\bigg{]}$
[…]
$\frac{1}{16L^{2}} {\mathbb{E}}\left[\frac{\left\|{{\boldsymbol{\varepsilon}}}_{t+1}\right\|^{2}}{\eta_{t}}\right]\leq\frac{1}{16L^{2}}{\mathbb{E}}\left[\frac{(1-\beta_{t+1})\left\|{{\boldsymbol{\varepsilon}}}_{t}\right\|^{2}}{\eta_{t}}\right]$
$~~+\frac{1}{8L^{2}}{\mathbb{E}}\left[\frac{\left\|\nabla f({\mathbf{x}}_{t+1},{\boldsymbol{\xi}}_{t+1})-\nabla f({\mathbf{x}}_{t},{\boldsymbol{\xi}}_{t+1})\right\|^{2}}{\eta_{t}}\right]$ (88)
Using $L$-smoothness, and the update equation (10), we obtain
$\frac{1}{16L^{2}}{\mathbb{E}}\left[\frac{\left\|{{\boldsymbol{\varepsilon}}}_{t+1}\right\|^{2}}{\eta_{t}}\right]$ $\leq \frac{1}{16L^{2}}{\mathbb{E}}\left[\frac{(1-\beta_{t+1})\left\|{{\boldsymbol{\varepsilon}}}_{t}\right\|^{2}}{\eta_{t}}\right]+\frac{1}{8}{\mathbb{E}}\left[\frac{\left\|{\mathbf{x}}_{t+1}-{\mathbf{x}}_{t}\right\|^{2}}{\eta_{t}}\right]$
$]\displaystyle~~~~~~~~~+\frac{1}{8L^{2}}{\mathbb{E}}\left[\frac{\beta_{t+1}^{2}G_{t+1}^{2}}{\eta_{t}}\right$ ] (90)
$\leq\frac{1}{16L^{2}}{\mathbb{E}}\left[\frac{(1-\beta_{t+1})\left\|{{\boldsymbol{\varepsilon}}}_{t}\right\|^{2}}{\eta_{t}}\right]+\frac{1}{8}{\mathbb{E}}\left[\eta_{t}\left\|{\boldsymbol{\delta}}_{t}\right\|^{2}\right]$
$~~~~~~~~~+\frac{1}{8L^{2}}{\mathbb{E}}\left[\frac{\beta_{t+1}^{2}G_{t+1}^{2}}{\eta_{t}}\right].$ (91)
Appendix D Proof of Lemma 2
Proof:
$\langle{{\hat{{\mathbf{x}}}}_{t}}-{\mathbf{x}}_{t},{\mathbf{z}}_{t+1}+{\mathbf{v}}_{{\hat{{\mathbf{x}}}}_{t}}\rangle$ (94)
$~+\langle{{\hat{{\mathbf{x}}}}_{t}}-{\mathbf{x}}_{t},\nabla{\tilde{f}}({{\hat{{\mathbf{x}}}}_{t}},{\mathbf{x}}_{t},{\mathbf{z}}_{t},{\boldsymbol{\xi}}_{t})-\nabla{\tilde{f}}({\mathbf{x}}_{t},{\mathbf{x}}_{t},{\mathbf{z}}_{t},{\boldsymbol{\xi}}_{t})\rangle\leq 0,$
$\Rightarrow\langle{\mathbf{z}}_{t+1}+{\mathbf{v}}_{{\hat{{\mathbf{x}}}}_{t}},{{\hat{{\mathbf{x}}}}_{t}}-{\mathbf{x}}_{t}\rangle+\mu\left\|{{\hat{{\mathbf{x}}}}_{t}}-{\mathbf{x}}_{t}\right\|^{2}\leq 0,$ (95)
Multiplying (95) by $\eta_{t}$ and adding, we obtain
Adding and subtracting $\nabla U({\mathbf{x}}_{t})$, we obtain
$\eta_{t}\langle{\mathbf{z}}_{t+1}-\nabla U({\mathbf{x}}_{t}),{{\hat{{\mathbf{x}}}}_{t}}-{\mathbf{x}}_{t}\rangle$
$~~~+\eta_{t}\langle\nabla U({\mathbf{x}}_{t}),{{\hat{{\mathbf{x}}}}_{t}}-{\mathbf{x}}_{t}\rangle$
$~~~\leq-\mu\eta_{t}\left\|{{\hat{{\mathbf{x}}}}_{t}}-{\mathbf{x}}_{t}\right\|^{2}+u({\mathbf{x}}_{t})-u({\mathbf{x}}_{t+1}).$ (98)
$\leq U({\mathbf{x}}_{t})+u({\mathbf{x}}_{t})-u({\mathbf{x}}_{t+1})-\eta_{t}\langle{\mathbf{z}}_{t+1}-\nabla U({\mathbf{x}}_{t}),{{\hat{{\mathbf{x}}}}_{t}}-{\mathbf{x}}_{t}\rangle$
$~~~+\bigg{(}\frac{L}{2}\eta_{t}^{2}-\mu\eta_{t}\bigg{)}\left\|{{\hat{{\mathbf{x}}}}_{t}}-{\mathbf{x}}_{t}\right\|^{2}$ (101)
${\mathbb{E}}\left[F({\mathbf{x}}_{t+1})-F({\mathbf{x}}_{t})\right]\leq\frac{1}{2}{\mathbb{E}}\left[\eta_{t}\left\|\nabla U({\mathbf{x}}_{t})-{\mathbf{z}}_{t+1}\right\|^{2}\right]$
$~~~+\frac{L}{2}{\mathbb{E}}\left[\eta_{t}^{2}\left\|{{\hat{{\mathbf{x}}}}_{t}}-{\mathbf{x}}_{t}\right\|^{2}\right]+\bigg{(}\frac{1}{2}-\mu\bigg{)}{\mathbb{E}}\left[\eta_{t}\left\|{{\hat{{\mathbf{x}}}}_{t}}-{\mathbf{x}}_{t}\right\|^{2}\right]$